OpenStudy (anonymous):

The height of an equilateral triangular prism increases by four units. The new volume is more than the original by how much?

OpenStudy (callisto):

area of equi. triangle = A original height = h original volume = 1/3 x A x h new vol. = 1/3 x A x 4h = 4 (1/3 x A x h ) = 4 (original volume ) % change = [ 4 (original volume ) - (original volume ) ] / (original volume ) *100% = + 300%

OpenStudy (anonymous):

Does that mean that the new volume is four more than the area of the base, more than the length of the height, times the length of the height, or times the area of the base?

OpenStudy (anonymous):

callisto's answer may be wrong.

OpenStudy (anonymous):

What would you do differently?

OpenStudy (callisto):

@fazna you're right, i overlooked something!

OpenStudy (callisto):

area of equi. triangle = A original height = h original volume = 1/3 x A x h new vol. = 1/3 x A x (h+4) = (1/3 x A x h ) + 4(1/3 x A) = (original volume ) + (4/3)(A) change in vol. =[ (original volume ) + (4/3)(A) ] - (original volume ) = (4/3)(A)

OpenStudy (anonymous):

Does that mean that the new volume is four more than the area of the base, more than the length of the height, times the length of the height, or times the area of the base?